# Dynamical Formation of Apparent Horizons due to Boundary Effect in Vacuum Einstein Gravity

### v2

@@ 5.5. Estimates on the Ricci coefficients
(122) $\begin{split}&\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\omega}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\\ \lesssim&\lVert(a^{\frac{1}{2}}\nabla)^{i}\underline{\omega}\rVert_{L^{2}_{(sc)}(S_{u,0})}+\int_{0}^{\underline{u}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\rho}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ +&\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}=i-1}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\rho}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ +&\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\psi_{g}\nabla^{i_{4}}\psi_{g}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ +&\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\underline{\omega}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ +&\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\underline{\omega}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}.\end{split}$
@@
(123) $\begin{split}&\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\underline{\omega}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ \lesssim&\int_{0}^{\underline{u}}a^{\frac{1}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\underline{\omega}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ \lesssim&\frac{a^{\frac{1}{2}}\Gamma^{2}}{\lvert u\rvert}\lesssim 1.\end{split}$
@@
(124) $\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\underline{\omega}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\frac{a\hskip 1.42262pt\Gamma^{3}}{\lvert u\rvert^{2}}\lesssim 1.$
@@
$\sum_{0\leq i\leq N+4}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\eta}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim\mathcal{R}[\beta]+1.$
@@
$\nabla_{4}\eta=\beta^{R}+(\hat{\chi},\mbox{tr}\chi)\cdot(\eta,\underline{\eta}),$
with the product of Yang-Mills curvature terms being once again gauge-invariant. Using the commutation formula 4.1 for the $\nabla_{4}$-direction, we have
@@
(126) $\begin{split}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\eta}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}&\lesssim\int_{0}^{\underline{u}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\beta}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ &+\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}=i-1}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\beta^{R}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ &+\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}(\eta,\underline{\eta})}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}.\end{split}$
For $0\leq i\leq N+4$, the first term is bounded by $\mathcal{R}[\beta]$. The second term is bounded by
(127) $\begin{split}\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}=i-1}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\beta}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\frac{\Gamma^{2}}{\lvert u\rvert}\lesssim 1.\end{split}$
@@
$\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}(\eta,\underline{\eta})}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\frac{\hskip 1.42262pt\Gamma^{2}}{\lvert u\rvert}\lesssim 1.$
@@
$\sum_{0\leq i\leq N+4}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\mbox{tr}\chi}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim\mathcal{R}[\alpha]+1.$
@@
Commuting with $i$ angular derivatives using 4.1, we obtain
@@
(129) $\begin{split}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\mbox{tr}\chi}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim&\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\hat{\chi}\nabla^{i_{4}}(\hat{\chi},\mbox{tr}\chi)}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ +&\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\psi_{g}\nabla^{i_{4}}\psi_{g}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ +&\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\psi_{g}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}.\end{split}$
@@
  • If in the term $\nabla^{i_{1}}\psi_{g}^{i_{2}}$ there exists some $\psi_{g}$ whose derivative is of order >N+3>N+3, we bound that term in $L^{2}_{(sc)}$ and the rest of the terms in $L^{\infty}_{(sc)}$. Notice, crucially, that from Remark LABEL:remarkhat, we can bound $\lVert{(a^{\frac{1}{2}}\nabla)^{i}\hat{\chi}}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})}\lesssim a^{\frac{1}{2}}$ for small $i$. As a consequence, we have the bound
    $\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\hat{\chi}\nabla^{i_{4}}(\hat{\chi},\mbox{tr}\chi)}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\frac{\lVert{(a^{\frac{1}{2}}\nabla)^{i}\hat{\chi}}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})}^{2}}{\lvert u\rvert}\lesssim 1.$
  • Otherwise, in the expression $\nabla^{i_{3}}\hat{\chi}\nabla^{i_{4}}\hat{\chi}$, at most one index $i_{3},i_{4}$ is greater than $N+1$ (in which case, we cannot bound that term in $L^{\infty}_{(sc)}$). Say without loss of generality, that $i_{3}>N+1$. We bound $(a^{\frac{1}{2}})^{i_{3}-1}\nabla^{i_{3}}\hat{\chi}$ in $L^{2}_{(sc)}$ above by $\mathcal{R}[\alpha]+1$ and the rest of the terms in $L^{\infty}_{(sc)}$ above by $1$ (using the improvement mentioned in Remark LABEL:remarkhat), whence
    $\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\hat{\chi}\nabla^{i_{4}}(\hat{\chi},\mbox{tr}\chi)}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\left(\mathcal{R}[\alpha]+1\right)\cdot 1.$
@@
Commuting with $i$ angular derivatives using 2, we get
@@
$\sum_{0\leq i\leq N+4}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\eta}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim\mathcal{R}[\beta]+\underline{\mathcal{R}}[\underline{\beta}]+1.$
@@
$\frac{1}{\lvert u\rvert}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\eta}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim\frac{\mathcal{R}[\beta]+\underline{\mathcal{R}}[\underline{\beta}]+1}{\lvert u\rvert}$,
@@ 6. $L^{2}(S_{u,\underline{u}})$ Estimates for the Weyl curvature components
$\sum_{0\leq i\leq N+2}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim 1.$
@@
Commuting with $i$ angular derivatives using 2, we arrive at
@@
(149) $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\lVert{a^{\frac{i}{2}}T_{i}^{4}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}+\lVert{a^{\frac{i}{2}}T_{i}^{5}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\right)\hskip 1.42262pt\text{d}u^{\prime}\lesssim\int_{u_{\infty}}^{u}\frac{a^{\frac{1}{2}}\Gamma^{2}}{\lvert u^{\prime}\rvert^{3}}\hskip 1.42262pt\text{d}u^{\prime}\lesssim 1.$
@@
(150) $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{a^{\frac{i}{2}}T_{i}^{6}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\int_{u_{\infty}}^{u}\frac{a^{\frac{1}{2}}\Gamma^{3}}{\lvert u^{\prime}\rvert^{2}}\hskip 1.42262pt\text{d}u^{\prime}\lesssim 1.$
@@
(151) $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{a^{\frac{i}{2}}T_{i}^{7}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\int_{u_{\infty}}^{u}\frac{a\Gamma^{2}}{\lvert u^{\prime}\rvert^{3}}\lesssim 1.$
@@
For $\Psi_{u}$ defined as in Section 4.6, there holds
$\sum_{0\leq i\leq N+3}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\Psi_{u}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim\mathcal{R}[\alpha]+1,$
@@
Commuting with $i$ angular derivatives using 4.1, we obtain
@@
$\hskip 19.91692pt\lVert{(a^{\frac{1}{2}}\nabla)^{i}\Psi_{u}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}$
$\lesssim$ $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\nabla^{i+1}\alpha}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}+\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\nabla^{i+1}\Psi_{u}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}$
@@
(153) $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\nabla^{i+1}\alpha}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\nabla^{i+1}\alpha}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{1/2}=\mathcal{R}[\alpha],$
@@
(154) $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\nabla^{i+1}\Psi_{u}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\frac{1}{a^{\frac{1}{2}}}\mathcal{R}\lesssim\frac{1}{a^{\frac{1}{2}}}R\lesssim 1,$

### v4

@@ 4.4. Estimates on the Ricci coefficients
(156) $\begin{split}&\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\omega}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\\ \lesssim&\lVert(a^{\frac{1}{2}}\nabla)^{i}\underline{\omega}\rVert_{L^{2}_{(sc)}(S_{u,0})}+\int_{0}^{\underline{u}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\rho}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ +&\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}=i-1}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\rho}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ +&\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\psi_{g}\nabla^{i_{4}}\underline{\omega}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ +&\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\underline{\omega}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ +&\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\underline{\omega}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}.\end{split}$
@@
(157) $\begin{split}&\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\underline{\omega}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ \lesssim&\int_{0}^{\underline{u}}a^{\frac{1}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\underline{\omega}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ \lesssim&\frac{\Gamma^{2}}{\lvert u\rvert^{2}}.\end{split}$
@@
(158) $\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\underline{\omega}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\frac{a\hskip 1.42262pt\Gamma^{3}}{\lvert u\rvert^{2}}.$
@@
$\sum_{0\leq i\leq N+4}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\eta}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim a^{-\frac{1}{2}}\mathcal{R}[\beta]+1.$
@@
$\nabla_{4}\eta=\beta+(\hat{\chi},\mbox{tr}\chi)\cdot(\eta,\underline{\eta}).$
Using the commutation formula 3.1 for the $\nabla_{4}$-direction, we have
@@
[…]
For $0\leq i\leq N+4$, the first term is bounded by $a^{-\frac{1}{2}}\mathcal{R}[\beta]$. The second term is bounded by
[…]
@@
$\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}(\eta,\underline{\eta})}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\frac{\hskip 1.42262pt\Gamma^{2}}{\lvert u\rvert}$.
@@
$\sum_{0\leq i\leq N+4}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\mbox{tr}\chi}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim\frac{a^{\frac{1}{2}}}{|u|}\mathcal{R}[\alpha]+1.$
@@
Commuting with $i$ angular derivatives using 3.1, we obtain
@@
[…]
@@
  • If in the term $\nabla^{i_{1}}\psi_{g}^{i_{2}}$ there exists some $\psi_{g}$ whose derivative is of order >N+3>N+3, we bound that term in $L^{2}_{(sc)}$ and the rest of the terms in $L^{\infty}_{(sc)}$. Notice, crucially, we can bound $\lVert{(a^{\frac{1}{2}}\nabla)^{i}\hat{\chi}}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})}\lesssim\frac{1}{a^{\frac{1}{2}}}$ for small $i$ in light of embedding theorems. As a consequence, we have the bound
    $\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\hat{\chi}\nabla^{i_{4}}(\hat{\chi},\mbox{tr}\chi)}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\frac{\lVert{(a^{\frac{1}{2}}\nabla)^{i}\hat{\chi}}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})}^{2}}{\lvert u\rvert}\lesssim\frac{1}{a|u|}.$
  • Otherwise, in the expression $\nabla^{i_{3}}\hat{\chi}\nabla^{i_{4}}\hat{\chi}$, at most one index $i_{3},i_{4}$ is greater than $N+1$ (in which case, we cannot bound that term in $L^{\infty}_{(sc)}$). Say without loss of generality, that $i_{3}>N+1$. We bound $(a^{\frac{1}{2}})^{i_{3}-1}\nabla^{i_{3}}\hat{\chi}$ in $L^{2}_{(sc)}$ above by $\mathcal{R}[\alpha]$ and the rest of the terms in $L^{\infty}_{(sc)}$ above by $1$, whence
    $\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\hat{\chi}\nabla^{i_{4}}(\hat{\chi},\mbox{tr}\chi)}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\left(\mathcal{R}[\alpha]+1\right)\cdot\frac{a^{\frac{1}{2}}}{|u|}.$
@@
Commuting with $i$ angular derivatives using 8, we get
@@
$\sum_{0\leq i\leq N+4}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\eta}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim a^{-\frac{1}{2}}\bigg(\mathcal{R}[\beta]+\underline{\mathcal{R}}[\underline{\beta}]\bigg)+1.$
@@
$\frac{1}{\lvert u\rvert}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\eta}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim\frac{a^{-\frac{1}{2}}(\mathcal{R}[\beta]+\underline{\mathcal{R}}[\underline{\beta}])+1}{\lvert u\rvert}$,
@@ 5. $L^{2}(S_{u,\underline{u}})$ Estimates for the Weyl curvature components
$\sum_{0\leq i\leq N+2}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim\frac{1}{a^{\frac{1}{2}}}.$
@@
Commuting with $i$ angular derivatives using 8, we arrive at
@@
(183) $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\lVert{a^{\frac{i}{2}}T_{i}^{4}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}+\lVert{a^{\frac{i}{2}}T_{i}^{5}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\right)\hskip 1.42262pt\text{d}u^{\prime}\lesssim\int_{u_{\infty}}^{u}\frac{a^{\frac{1}{2}}\Gamma^{2}}{\lvert u^{\prime}\rvert^{3}}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{a^{\frac{1}{2}}\Gamma^{2}}{|u|^{2}}.$
@@
(184) $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{a^{\frac{i}{2}}T_{i}^{6}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\int_{u_{\infty}}^{u}\frac{a^{\frac{1}{2}}\Gamma^{3}}{\lvert u^{\prime}\rvert^{2}}\frac{a^{\frac{1}{2}}}{|u^{{}^{\prime}}|}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{a\Gamma^{3}}{|u|^{2}}\lesssim\frac{a^{\frac{1}{2}}}{|u|}.$
@@
(185) $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{a^{\frac{i}{2}}T_{i}^{7}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\int_{u_{\infty}}^{u}\frac{a\Gamma^{2}}{\lvert u^{\prime}\rvert^{3}}\frac{a^{\frac{1}{2}}}{|u^{{}^{\prime}}|}\lesssim\frac{a^{\frac{3}{2}}\Gamma^{2}}{|u|^{3}}.$
@@
For $\Psi_{u}=(\beta,\underline{\beta},\rho,\sigma)$ defined as in Section 3.6, there holds
$\sum_{0\leq i\leq N+3}\lVert{(a^{\frac{1}{2}}\nabla)^{i}(\beta,\underline{\beta})}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim a^{-\frac{1}{2}}(\mathcal{R}[\alpha]+1),$
@@
Commuting with $i$ angular derivatives using 3.1, we obtain
@@
[…]
$+\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\nabla^{i+1}\alpha}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}+\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\nabla^{i+1}\Psi_{u}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}$
@@
(187) $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\nabla^{i+1}\alpha}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\nabla^{i+1}\alpha}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{1/2}=\frac{1}{a}\mathcal{R}[\alpha],$
@@
(188) $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\nabla^{i+1}\Psi_{u}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\frac{1}{a^{\frac{1}{2}}}\mathcal{R}+\frac{1}{a}\mathcal{R}[\alpha],$